OpenStudy (anonymous):

Which of the following is the rational exponent expression of third root of 4 n? (4n)3 3n4 quantity of 4 n to the one third power 4 times n to the one third power

OpenStudy (anonymous):

\[\sqrt[3]{4n}\]

OpenStudy (anonymous):

last option?

OpenStudy (anonymous):

\[4n1/3\]

OpenStudy (anonymous):

for this you have to know the rule of radicals notation \[\sqrt[7]{2^3}=2^{\frac{3}{7}}\]

OpenStudy (anonymous):

I believe it is the second last because you have to take the root of 4n and not take the root of 4 only

OpenStudy (anonymous):

of n only*

OpenStudy (anonymous):

so (4n)1/3 would only allow us to take the roof of n?

OpenStudy (anonymous):

(4n)^1/3 takes root of (4n), it is correct "4 times n to the one third power" : 4 * (n)^1/3

OpenStudy (anonymous):

alright I gotcha

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

quantity of 4n to the one third power